OpenStudy (anonymous):

An airplane flies at airspeed (speed relative to the air) 187 km/h. The pilot wishes to fly due West but there is a 57.7 km/h wind blowing in from North to South. In what direction should the pilot head the plane? Answer in units of  (north of West)

OpenStudy (anonymous):

What would be the ground speed of the plane (its speed relative to the ground)? Answer in units of km/h

OpenStudy (anonymous):

\[v_{w} = (57.7km/hr)\]\[v_{t} = (187km/hr)\] \[v_{a} = v_{t} cos (theta)\] \[v_{w} = v_{t} sin (theta)\]

OpenStudy (anonymous):

It's just a 187, 57.5 right triangle Your new \[\theta\] is just 57.5/187sin^-1\[\theta\] =17.9 degrees above due west (17.9 degrees west-northwest)

OpenStudy (anonymous):

(PS, I can't read the units you require)

OpenStudy (anonymous):

Ground speed is \[\sqrt{187^{2}-57.5^{2}}\] = 177.94 kmh